OpenStudy (anonymous):

solve |x/3 + 2/5| =2

hero (hero):

Hint: x/3 + 2/5 = 2 x/3 + 2/5 = -2

OpenStudy (anonymous):

why is it positive AND negative 2? doesn't absolute value make it positive only? @Hero

hero (hero):

I'll explain by demonstration: |2| = 2 |-2| = 2

hero (hero):

Notice that with absolute value, there are two values of which the absolute value is 2

hero (hero):

So it makes sense that there are two equations that can represent this situation.

OpenStudy (anonymous):

oh okay, i forgot about that rule, thanks! so i did both of the equations you hinted and got 24/5 and -36/5 :O

OpenStudy (anonymous):

ill type out my work, just give me a minute :)

hero (hero):

Typing out your work is always the best thing to do.

OpenStudy (anonymous):

\[\frac{5}{5}*\frac{x}{3}+\frac{2}{5}*\frac{3}{3}=2\] \[\frac{5(x)}{15}+\frac{2(3)}{15}=2\] \[\frac{5x}{15}+\frac{6}{15}=2\] \[\frac{15}{1}*\frac{5x+6}{15}=2*\frac{15}{1}\] \[5x+6=30\] \[5x=24\] \[x=\frac{24}{5}\]

hero (hero):

Yes, that's correct.

OpenStudy (anonymous):

\[\frac{5}{5}*\frac{x}{3}+\frac{2}{5}*\frac{3}{3}=-2\] \[\frac{5x}{15}+\frac{6}{15}=-2\] \[5x+6=-30\] \[5x=-36\] \[x=\frac{-36}{5}\]

OpenStudy (anonymous):

thank you so much for checking and helping me! i appreciate it :)

hero (hero):

You can always check your work: Notice that: \[\frac{\frac{24}{5}}{3} + \frac{2}{5}\] \[\frac{24}{5}\frac{1}{3} + \frac{2}{5}\] \[\frac{24}{15} + \frac{6}{15} = \frac{30}{15} = 2\]

OpenStudy (anonymous):

oh wow, i didn't know that was how to check the work! thanks that really helps! :)

hero (hero):

Of course if you used -36/5 in place of x, you would end up with -2. But then the absolute value of -2 is 2 so it checks out :)